A baseball of mass 0.15 kg is thrown at a batter at a speed of 42 m/s, and, after being hit, travels in the opposite direction.

If the final speed of the baseball is 35 m/s and the baseball and the bat are in contact for 2 ms, what is the average force the bat applies to the baseball?

What percentage of the original kinetic energy is lost?
Answer #1

for the average force

average force on the ball = change in momentum/time taken

average force on the ball = m * (v - u)/time

average force on the ball = 0.15 * (35 - (-42))/0.002

average force on the ball = 5775 N

the average force on the ball is 5775 N

----

percentage of original kinetic energy lost = (-0.50 *m * v^2 + 0.50 *m * u^2)/(0.50 * m * u^2)

percentage of original kinetic energy lost = (42^2 - 35^2)/42^2

percentage of original kinetic energy lost = 0.306 = 30.6 %

the percentage of original kinetic energy lost is 30.6 %